OpenStudy (anonymous):

Solve for x:−3(x + 3) = −3(x + 1) − 5

mathslover (mathslover):

first of all \[\huge{\textbf{Welcome to openstudy}}\] given : \[\huge{-3(x+3)=-3(x+1)-5}\] \[\huge{-3(x)-3(3)=-3(x)-3(1)-5}\] \[\huge{-3x-9=-3x-3-5}\] \[\huge{-3x+3x=-3-5+9}\] \[\huge{0=1}\] that is not true hence no solutions for the given equation